# Corruption-Robust Linear Bandits: Minimax Optimality and Gap-Dependent Misspecification

### v1

@@ Appendix H Computationally Efficient Algorithm for Adversarial $C$ Bound
H.1 Preliminaries for Entropic Barrier
@@
H.2 Auxiliary Lemmas
Lemma H.1 (Lemma 1 of Ito et al., (2020)).
@@ H.2 Auxiliary Lemmas
Lemma H.3 (Lemma 14 of Zimmert and Lattimore, (2022)).
@@
Lemma H.4 (Proposition 2.3.2 in Nesterov and Nemirovskii, (1994)).
@@ H.3 Regret Analysis
The individual terms Bias, Deviation, Bonus, Penalty, Stability terms are bounded in Lemma H.6, Lemma H.7, Lemma H.9, Lemma H.10, Lemma H.12.
Lemma H.5.
@@
Define $y=\Sigma_{t}^{-\frac{1}{2}}a$. Then $\mathbb{E}_{y}\left[yy^{\top}\right]=\Sigma_{t}^{-\frac{1}{2}}\mathbb{E}_{a\sim p_{t}}[aa^{\top}]\Sigma_{t}^{-\frac{1}{2}}=I$. Since $p_{t}$ is a log-concave distribution, and log-concavity is preserved under liner transformation, $y$ is also log-concave. Applying Lemma H.1 on it leads to
@@
The proof is the same as that of Lemma LABEL:lem:_bias_term_cew. ∎
Lemma H.7.
@@
The proof is the same as that of Lemma LABEL:lem:_dev_cew. ∎
Lemma H.8.
@@
$\leq-\frac{\alpha}{\gamma^{\prime}}\sum_{\tau\in\mathcal{I}}\|u-x_{\tau}\|_{\nabla^{2}F(x_{\tau})}+4\alpha d|\mathcal{I}|+\frac{2\alpha\sqrt{d}|\mathcal{I}|}{\gamma^{\prime}}$ (Lemma H.3 and $F$ is $d$-self-concordant barrier)
$\leq-\frac{\alpha}{2\gamma^{\prime}}\sum_{\tau\in\mathcal{I}}\|u-x_{\tau}\|_{\widetilde{\Sigma}_{\tau}^{-1}}+\mathcal{O}(\alpha d^{2}\log T)$ ($\nabla^{2}F(x_{\tau})=\Sigma_{\tau}^{-1}\succeq\frac{1}{4}\widetilde{\Sigma}_{\tau}^{-1}$ and Lemma H.8)
@@
Since $x^{\star}=\min_{x\in{\mathcal{A}}}F(x)$ and $\pi_{x^{\star}}(u)\leq 1-\frac{1}{T}$ from Eq. (23). We have from Lemma H.4
@@
Lemma H.11 (Lemma 17 in Zimmert and Lattimore, (2022)).
@@
By Lemma H.2, we have $\widetilde{\Sigma}_{t}^{-1}\preceq\left(\mathbb{E}_{a\sim\tilde{p}_{t}}[aa^{\top}]\right)^{-1}\preceq 2\Sigma_{t}^{-1}$, and thus
@@
Thus, $\|\widehat{\theta}_{t}+b_{t}\|_{\nabla^{2}F(x_{t})^{-1}}\leq\beta\sqrt{2d}+\alpha\sqrt{d}$. If $\eta\leq\frac{1}{16(\beta\sqrt{2d}+\alpha\sqrt{d})}$, by Lemma H.11, we have
@@
Putting Lemma H.6, Lemma H.7, Lemma H.9, Lemma H.10, Lemma H.12 into Eq. (24) and Eq. (25), with $\eta\leq\frac{1}{16(\beta\sqrt{2d}+\alpha\sqrt{d})}$ and $\gamma=\frac{1}{T}$, we have with probability at least $1-\mathcal{O}(\delta)$,
@@
Appendix I Gap-dependent Misspecification
@@ 6       Calculate
Define $\text{\rm Gap}(x)=f_{0}^{\star}-f_{0}(x)$ as the suboptimal gap at point $x$. Definition 1 implies the true value function $f_{0}(x)=\theta^{\top}x+\Delta(x)$ where $|\Delta(x)|\leq\rho(f_{0}^{\star}-f_{0}(x))=\rho\text{\rm Gap}(x)$. We further assume that $|\Delta(x)|\leq\rho\text{\rm Gap}(x)$ which captures both standard uniform misspecification and the gap-dependent misspecification. With this assumption, our main result is summarized in Theorem I.1.
Theorem I.1.
@@
When $|{\mathcal{A}}|\geq T^{d}$, we can apply similar covering number arguments as in Appendix E, replacing ${\mathcal{A}}_{1}$ with a $\frac{1}{T}$-net of ${\mathcal{A}}$. Combined with Theorem I.1, this yields the result in Theorem 6.2.
@@
Appendix J General Reduction from Corruption-Robust Algorithms to Misspecification
@@
Under Assumption 4, we have that $\mathbb{E}^{\mathcal{M}^{\star},\pi^{\star}}\left[\sum_{h=1}^{H}\epsilon^{\mathrm{mis}}_{h}(s_{h},a_{h})\right]=0$. Lemma J.2 then implies that
@@
Furthermore, Lemma J.2 also implies
@@
Next, note that Theorem 4.2 of Ye et al., (2023) gives an algorithm on $\mathcal{M}_{0}$ satisfying Assumption 5 with $\mathcal{C}_{1}=\widetilde{\mathcal{O}}(\sqrt{H^{2}d^{3}})$ and $\mathcal{C}_{2}=\widetilde{\mathcal{O}}(Hd)$ (assuming that $\sum_{h=1}^{H}r_{h}\in[0,1]$ almost surely). We can then apply Theorem J.1 to obtain the result.
Appendix K Auxiliary Lemmas
Lemma K.1 (Lemma 16 of Zimmert and Lattimore, (2022)).
@@
Lemma K.3 (Strengthened Freedman’s inequality (Theorem 9 of Zimmert and Lattimore, (2022))).

### v2

@@ 3       Define for $y\in\phi({\mathcal{X}})$,
J.1 Preliminaries for Entropic Barrier
@@
J.2 Auxiliary Lemmas
Lemma J.1 (Lemma 1 of Ito et al., (2020)).
@@
Lemma J.3 (Lemma 14 of Zimmert and Lattimore, (2022)).
@@
Lemma J.4 (Proposition 2.3.2 in Nesterov and Nemirovskii, (1994)).
@@
The individual terms Bias, Deviation, Bonus, Penalty, Stability terms are bounded in Lemma J.6, Lemma J.7, Lemma J.9, Lemma J.10, Lemma J.12.
Lemma J.5.
@@
Define $y=\Sigma_{t}^{-\frac{1}{2}}a$. Then $\mathbb{E}_{y}\left[yy^{\top}\right]=\Sigma_{t}^{-\frac{1}{2}}\mathbb{E}_{a\sim p_{t}}[aa^{\top}]\Sigma_{t}^{-\frac{1}{2}}=I$. Since $p_{t}$ is a log-concave distribution, and log-concavity is preserved under liner transformation, $y$ is also log-concave. Applying Lemma J.1 on it leads to
@@
The proof is the same as that of Lemma H.1. ∎
Lemma J.7.
@@
The proof is the same as that of Lemma H.2. ∎
Lemma J.8.
@@
$\leq-\frac{\alpha}{\gamma^{\prime}}\sum_{\tau\in\mathcal{I}}\|u-x_{\tau}\|_{\nabla^{2}F(x_{\tau})}+4\alpha d|\mathcal{I}|+\frac{2\alpha\sqrt{d}|\mathcal{I}|}{\gamma^{\prime}}$ (Lemma J.3 and $F$ is $d$-self-concordant barrier)
$\leq-\frac{\alpha}{2\gamma^{\prime}}\sum_{\tau\in\mathcal{I}}\|u-x_{\tau}\|_{\widetilde{\Sigma}_{\tau}^{-1}}+\mathcal{O}(\alpha d^{2}\log T)$ ($\nabla^{2}F(x_{\tau})=\Sigma_{\tau}^{-1}\succeq\frac{1}{4}\widetilde{\Sigma}_{\tau}^{-1}$ and Lemma J.8)
@@
Since $x^{\star}=\min_{x\in{\mathcal{A}}}F(x)$ and $\pi_{x^{\star}}(u)\leq 1-\frac{1}{T}$ from Eq. (25). We have from Lemma J.4
@@
Lemma J.11 (Lemma 17 in Zimmert and Lattimore, (2022)).
@@
By Lemma J.2, we have $\widetilde{\Sigma}_{t}^{-1}\preceq\left(\mathbb{E}_{a\sim\tilde{p}_{t}}[aa^{\top}]\right)^{-1}\preceq 2\Sigma_{t}^{-1}$, and thus
@@
Thus, $\|\widehat{\theta}_{t}+b_{t}\|_{\nabla^{2}F(x_{t})^{-1}}\leq\beta\sqrt{2d}+\alpha\sqrt{d}$. If $\eta\leq\frac{1}{16(\beta\sqrt{2d}+\alpha\sqrt{d})}$, by Lemma J.11, we have
@@
Putting Lemma J.6, Lemma J.7, Lemma J.9, Lemma J.10, Lemma J.12 into Eq. (26) and Eq. (27), with $\eta\leq\frac{1}{16(\beta\sqrt{2d}+\alpha\sqrt{d})}$ and $\gamma=\frac{1}{T}$, we have with probability at least $1-\mathcal{O}(\delta)$,
@@
Appendix K Gap-dependent Misspecification
@@ 6       Calculate
Define $\text{\rm Gap}(x)=f_{0}^{\star}-f_{0}(x)$ as the suboptimal gap at point $x$. Definition 3 implies the true value function $f_{0}(x)=\theta^{\top}x+\Delta(x)$ where $|\Delta(x)|\leq\rho(f_{0}^{\star}-f_{0}(x))=\rho\text{\rm Gap}(x)$. We further assume that $|\Delta(x)|\leq\rho\text{\rm Gap}(x)$ which captures both standard uniform misspecification and the gap-dependent misspecification. With this assumption, our main result is summarized in Theorem K.1.
Theorem K.1.
@@
When $|{\mathcal{A}}|\geq T^{d}$, we can apply similar covering number arguments as in Appendix E, replacing ${\mathcal{A}}_{1}$ with a $\frac{1}{T}$-net of ${\mathcal{A}}$. Combined with Theorem K.1, this yields the result in Theorem 6.2.
@@
Appendix L General Reduction from Corruption-Robust Algorithms to Misspecification
@@
Under Assumption 4, we have that $\mathbb{E}^{\mathcal{M}^{\star},\pi^{\star}}\left[\sum_{h=1}^{H}\epsilon^{\mathrm{mis}}_{h}(s_{h},a_{h})\right]=0$. Lemma L.2 then implies that
@@
Furthermore, Lemma L.2 also implies
@@
Next, note that Theorem 4.2 of Ye et al., (2023) gives an algorithm on $\mathcal{M}_{0}$ satisfying Assumption 5 with $\mathcal{C}_{1}=\widetilde{\mathcal{O}}(\sqrt{H^{2}d^{3}})$ and $\mathcal{C}_{2}=\widetilde{\mathcal{O}}(Hd)$ (assuming that $\sum_{h=1}^{H}r_{h}\in[0,1]$ almost surely). We can then apply Theorem L.1 to obtain the result.
Appendix M Auxiliary Lemmas
Lemma M.1 (Lemma 16 of Zimmert and Lattimore, (2022)).
@@
Lemma M.3 (Strengthened Freedman’s inequality (Theorem 9 of Zimmert and Lattimore, (2022))).